OpenStudy (anonymous):

Morissa pays $12.99 per pound for dried blueberries and $9.99 per pound for dried cherries. Which of the following statements could be true? Morissa bought 2 pounds of blueberries and spent $14.99 on cherries for a total of $34.99. Morissa bought 1.3 pounds of blueberries and 1.6 pounds of cherries for a total of $33.77. Morissa spent $22.73 on blueberries and spent $12.49 on cherries for a total of 3 pounds of fruit. Morissa bought 2 pounds of fruit for a total of $18.

OpenStudy (anonymous):

@heather040200 @primeralph

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

So lets do the math... Option A: 2 pounds of blueberries @ 12.99 per pound = 25.98 + 14.99 spent on cherries = 40.97 not 34.99 Option B: 1.3 pounds of blueberries @ 12.99 per pound, so 1.3*12.99=16.89 1.6 pounds of cherries @ 9.99 per pound, so 1.6*9.99=15.98 Finally add both results 16.89+15.98=32.87, not 33.77 Option C: blueberries- 22.73 dollars spent divided by 12.99 price per pound, so 22.73/12.99=1.75lbs of blueberries cherries- 12.49 dollars spent divided by 9.99 price per pound, so 12.49/9.99=1.25 Now add both weight results together 1.75+1.25=3 Answer is Option C: TRUE- Morissa spent $22.73 on blueberries and spent $12.49 on cherries for a total of 3 pounds of fruit